a baseball pitcher throws the ball. If the linear velocity of the ball relative to the elbow joint 23.0 (m)/(s) at a distance of 0.340m from the joint and the moment of inertia of the forearm is 0.5 kg*m^(2), what is the rotational kinetic energy?


1
Expert's answer
2021-01-16T17:24:11-0500

By definition, the rotational kinetic energy is:


"K = \dfrac{I\omega^2}{2}"

where "I = 0.5 kg\cdot m^2" is the moment of inertia, and "\omega" is the angular frequency. The angular frequency relates to the linear velocity as follows:


"\omega = \dfrac{v}{R}"

where "v = 23m\/s", and "R = 0.34m" is the distance from the joint. Substituting "\omega" into "K", obtain:


"K = \dfrac{Iv^2}{2R^2}\\
K = \dfrac{0.5\cdot 23^2}{2\cdot 0.34^2} \approx 1144J"

Answer. 1144 J.